OpenStudy (anonymous):

(4x)/(3(1x*2))*(x^2-4)/(x)

OpenStudy (anonymous):

must make x^2-4 and perfect square. I am unsure how to do so.

OpenStudy (anonymous):

Google crashed.. here we write it again :(

OpenStudy (anonymous):

\[\LARGE {4x \over 3(1x\cdot 2)}\cdot {x^2-4 \over x}=\] \[\LARGE {4x \over 3x\cdot 6}\cdot {x^2-4 \over x}=\] \[\LARGE {4\cancel{x} \over 18\cancel{x}}\cdot {x^2-4 \over x}=\] \[\LARGE {2 \over 9}\cdot {x^2-4 \over x}=\] \[\LARGE {2(x^2-4) \over 9x}\]

OpenStudy (anonymous):

is this what you're asking for.. ?

OpenStudy (anonymous):

I think.. The original problem was (4x)/(3x+6)*(x^2-4)/(x) and i factored the 3x + 6 part.. And i need to find the perfect square of the top right equation.. i am unsure how to make that section a perfect square and then simplify the expression..

OpenStudy (anonymous):